OpenStudy (anonymous):

i do not understand how to do any of this its quadratic functions can someone help

OpenStudy (anonymous):

where are the quadratic functions?

OpenStudy (anonymous):

its asking me to find the equation of the axis of symmetry the problem is y=x^2-3

OpenStudy (mrnood):

do oyu understand that the equation can be written in 'general' form as \[y= ax ^{2}+bx + c\] and its roots are when y=0

OpenStudy (mrnood):

if you look at YOUR equation can you write the values of a,b and c?

OpenStudy (mrnood):

I cannot help because you went offline 2 minutes after posting your question.

OpenStudy (mrnood):

If oyu can identify the values of a b and c in your equation based on the standard form I gave above, then I will help you find the axis of symmetry

OpenStudy (anonymous):

i dont have wifi at my house and i left school is why i went offlline

OpenStudy (anonymous):

and i dont know if i can find the values of a and b. it doenst tell me to do that it just tells me to find the axis of symmetry

OpenStudy (mrnood):

you need to do what I asked to answer the question - it is not 'extra' work it is part of th eprocess to solve oyur question. In GENERAL (that is for ALL quadratic equations) you can writhe them in the form ax^2 + bx + c =0 for example if the equation is 4x^2 -3x +27 then a = 4 b= -3 c= 27 YOUR equation is y= x^2 - 3 so compare your equation to my example and see if you can find the equivalent values for a, b and c

OpenStudy (anonymous):

is this right? a=y b=x^2 c=3

OpenStudy (mrnood):

no sorry a is the number in front of the x^2 b is the number in front of the x and c is th econstant term, with no x in it Have another look and try again..

OpenStudy (anonymous):

how do i figure out which is which?

OpenStudy (mrnood):

your equation is y = x^2 - 3 you need the form y = a(x^2) +bx + c just look at the general form, and compare it with your equation find a value for a so that ax^2 is the same as x^2 find a value for b so that bx = 0 find a value for c such that c= -3 If you are having trouble doing this then I STRONGLY recommend that a) you review your text, b) look at this: http://en.wikipedia.org/wiki/Quadratic_equation c) and talk to your teacher. This is all fundamental stuff about quadratics and parabolas You must get an understanding of the form ax^2 + bx +c = 0 before you can do much more in this subject